# Almost equivalences between Tamarkin category and Novikov sheaves

### v2

@@ 7.3 Relation to non-conic microsupport
For an object $\mathcal{E}\in\mu^{\mathbb{G}}(T^{*}M)$, we have
@@
Let $(x,\xi)$ be a point in $T^{*}M$. Consider any $C^{1}$-function with $\phi$ with $d\phi(x)=\xi$. For any $c^{\prime}>0$, the equivariant sheaf $\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{-\phi+c+c^{\prime}>t\geq-\phi+c}$ is almost sent to $R_{c}^{\phi}$ on sufficiently small open subset around $x$ under $\mathfrak{A}$. Hence the test sheaves to estimate the both sides of the desired equality coincide. This completes the proof. ∎
@@ 8 Application II: Curved sheaves
In Fukaya category theory, we have to deal with curved complexes, bounding cochain, bulk deformations. In sheaf theory (in the setup of Tamarkin category), introducing such notions is not easy (although we can do them as partly explained in [IK]). Our interpretation of the category $\mathrm{Sh}^{\mathbb{R}}_{\tau>0}(X\times\mathbb{R}_{t})$ as the sheaf category of $\Lambda_{0}$-modules allows us to introduce such deformations easily.
@@ 8.1 Curved complex and sheaves
We denote the full subcategory spanned by acyclic complexes by $Acycl$, We set the Drinfeld quotient by
@@
Associated to $\mathrm{CSh}(X,R)$, the derived dg category of curved sheaves.
@@ 8.2 Twisted sheaves
For the sake of simplicity, we assume $\mathbb{K}=\mathbb{C}$.
@@
For general coefficients, the same proof wokrs by replacing the de Rham resolution with Cech resolution. ∎

### v3

@@ 7.3 Relation to non-conic microsupport
For an object $\mathcal{E}\in\mu^{\mathbb{G}}(T^{*}X)$, we have
@@
Let $(x,\xi)$ be a point in $T^{*}X$. Consider any $C^{1}$-function with $\phi$ with $d\phi(x)=\xi$. For any $c^{\prime}>0$, the equivariant sheaf $\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{-\phi+c+c^{\prime}>t\geq-\phi+c}$ is almost sent to $R_{c^{\prime}}^{\phi}$ on sufficiently small open subset around $x$ under $\mathfrak{A}$. Hence the test sheaves to estimate the both sides of the desired equality coincide. This completes the proof. ∎
@@ 8 Application II: Curved sheaves
In Fukaya category theory, we have to deal with curved complexes, bounding cochains, and bulk deformations. In sheaf theory (in the setup of Tamarkin category), introducing such notions is not easy (although we can do them as partly explained in [IK]). Our interpretation of the category $\mathrm{Sh}^{\mathbb{R}}_{\tau>0}(X\times\mathbb{R}_{t})$ as the sheaf category of $\Lambda_{0}$-modules allows us to introduce such deformations easily.
@@ 8.1 Curved complex and sheaves
We denote the full subcategory spanned by the totalizations of the acyclic complexes by $Acycl$, We set the Drinfeld quotient by
@@
This is the derived dg category of curved sheaves.
@@ 8.2 Twisted sheaves
For the sake of simplicity, we assume $\mathbb{K}=\mathbb{C}$. Then we can use the de Rham model for the cohomology e.g. $H^{2}(X,\mathfrak{m})\subset\prod_{c\in\mathbb{R}_{>0}}H^{2}(X,\mathbb{C})\cong\prod_{c\in\mathbb{R}_{>0}}H^{2}_{dR}(X,\mathbb{C})$.
@@
For general coefficients, the same proof works by replacing the de Rham resolution with Cech resolution. ∎